# Exploring the Most Extreme Gamma-Ray Blazars Using Broadband Spectral Energy Distributions

### v1

@@ Abstract
Methods. A sample of 124 $\gamma$-ray blazars is analysed, selected for their high synchrotron peak frequencies and $\gamma$-ray emission properties, with a focus on sources showing low variability and good broadband data coverage. Their SEDs are constructed using archival multi-wavelength data from the SSDC SED Builder service, supplemented with recent Swift-UVOT, Swift-XRT, and Fermi-LAT observations. The SEDs are modelled with a one-zone synchrotron/synchrotron-self-Compton framework, classifying sources by synchrotron peak frequency. EHSP properties are compared to other blazar populations, and their detectability with the Cherenkov Telescope Array Observatory (CTAO) is assessed.
Results. We identify 66 new EHSP candidates, significantly expanding the known population. Their synchrotron peak frequencies are statistically higher than in previous studies, and they exhibit low Compton dominance, consistent with environments lacking strong external photon fields. A clear correlation between synchrotron peak frequency and the magnetic-to-kinetic energy density ratio is found, with the most extreme EHSPs nearing equipartition. Host galaxy emission is detected in many sources, but no significant differences are observed between elliptical and lenticular hosts. Finally, our analysis suggests that 9 high-synchrotron peaked/EHSPs could be observed by CTAO at $>5\sigma$ (20 at $>3\sigma$) in 20-hour exposures, a feasible integration time for Imaging Atmospheric Cherenkov Telescopes, highlighting their potential to improve studies of extreme jet physics and cosmology.
@@ 1 Introduction
Active galactic nuclei (AGNs) are extreme cosmic sources, powered by matter accretion onto a supermassive black hole (SMBH, e.g., Combes, 2021). Some develop relativistic jets extending to kiloparsec scales (e.g., Marschern, 2010). Unification models classify radio-loud AGNs by jet viewing angle (e.g., Urry & Padovani, 1995), with blazars, jets aligned close to our line of sight, dominating the extragalactic $\gamma$-ray sky (Abdollahi et al., 2020).
@@
Blazars have broadband spectral energy distributions (SEDs) dominated by non-thermal emission from radio to $\gamma$ rays (e.g., Fossati et al., 1998; Padovani, 2017). Their SEDs feature a double-peaked structure: the lower-energy peak (infrared to X-ray) arises from synchrotron emission by relativistic jet electrons, while the origin of the higher-energy $\gamma$-ray peak remains debated (e.g., Abdo et al., 2010), with both hadronic and leptonic processes proposed.
@@
Leptonic models explain the higher-energy peak of blazar SEDs as inverse Compton (IC) scattering, where high-energy electrons interact with low-energy photons (e.g., MAGIC Collaboration et al., 2020, 2021). This occurs mainly through synchrotron-self-Compton (SSC), where electrons scatter their own synchrotron photons (e.g., Maraschi et al., 1992; Finke et al., 2008), and external Compton, where electrons scatter external photons from sources like the cosmic microwave background, broad-line region (BLR), dusty torus, disk, stars, or different jet layers (e.g., Sikora et al., 1994; van den Berg et al., 2019). Hadronic models instead attribute $\gamma$-ray emission to high-energy proton interactions with ambient matter or radiation fields and to synchrotron radiation from relativistic protons in strong magnetic fields (e.g., Mannheim & Biermann, 1992; Böttcher et al., 2013, and references therein). Lepto-hadronic models combine both leptonic and hadronic processes (e.g., Mücke et al., 2003; Paliya et al., 2018). The main acceleration mechanisms proposed for ultra-relativistic particles in blazar jets include shock acceleration via repeated shock crossings and magnetic turbulence (e.g., Kirk et al., 1998; Tavecchio et al., 2020), as well as magnetic reconnection (e.g., Sironi & Spitkovsky, 2014) and stochastic acceleration (e.g., Blandford & Eichler, 1987; Liu et al., 2017).
@@
Blazars are classified based on their optical spectra into flat spectrum radio quasars (FSRQs) and BL Lacertae objects (BL Lacs). FSRQs exhibit strong, broad emission lines (EW $>5$), while BL Lacs have an almost featureless continuum (e.g., Marcha et al., 1996), indicating that external photon sources like the BLR, dusty torus, or disk are subdominant, with emission mainly from synchrotron/SSC (e.g., Ghisellini et al., 2011). In the most extreme BL Lacs, where the synchrotron peak is at higher frequencies, the host galaxy’s thermal emission becomes visible in the optical/ultraviolet (UV) range, allowing absorption lines to be detected. Blazars are also categorized by their synchrotron peak frequency (e.g., Padovani & Giommi, 1995; Abdo et al., 2010), from low-synchrotron peaked (LSP, $\nu_{sync}^{peak}<10^{14}$ Hz) to intermediate (ISP, $10^{14}$ Hz $\leq\nu_{sync}^{peak}<10^{15}$ Hz), high (HSP, $10^{15}$ Hz $\leq\nu_{sync}^{peak}<10^{17}$ Hz), and extreme high-synchrotron peaked (EHSP, $\nu_{sync}^{peak}\geq 10^{17}$ Hz), with peaks often reaching the X-ray band.
Blazars often show variability on timescales from minutes to years, differing across energy bands (e.g., Urry, 1996; Marscher et al., 2008). HSPs and EHSPs vary significantly in X-rays but show little to no large-scale variability in the $\gamma$-ray band, whereas LSPs and ISPs are generally more variable at high-energy (HE, 100 MeV $<E<$ 100 GeV) $\gamma$ rays. This pattern may result from the low $\gamma$-ray luminosity of HSPs and EHSPs, as their apparent lack of variability could stem from the limited sensitivity of Fermi-LAT and Imaging Atmospheric Cherenkov Telescopes (IACTs) to low-flux sources (Ackermann et al., 2011).
@@
FSRQs, the most luminous blazars, are mainly LSPs, while higher synchrotron peak frequency blazars are mostly lower-luminosity BL Lacs (e.g., Giommi et al., 2012). This anti-correlation between luminosity and synchrotron peak frequency defines the blazar sequence (Fossati et al., 1998; Ghisellini et al., 2017), though observational biases and sample selection may challenge its validity (e.g., Padovani, 2007; Giommi et al., 2012; Padovani et al., 2022).
@@
EHSPs, with radiatively inefficient accretion disks, occupy the end of the blazar sequence (e.g., Ghisellini et al., 2011). Their extreme synchrotron peak frequencies challenge standard models, requiring extreme conditions for the ultra-relativistic particles responsible for their emission (e.g., Biteau et al., 2020). Studying EHSPs enhances our understanding of jet acceleration and non-thermal emission. Their high energies (up to tens of TeV, Biteau et al., 2020) make them valuable for cosmological studies, including the extragalactic background light (EBL, e.g., Domínguez & Ajello, 2015; Domínguez et al., 2019; Gréaux et al., 2024), blazar evolution (e.g., Ajello et al., 2014), the extragalactic $\gamma$-ray background (e.g., Ajello et al., 2015; Paliya et al., 2019), and the intergalactic magnetic field (e.g., Bonnoli et al., 2015). As potential very-high-energy (VHE, $E>100$ GeV) emitters, they are prime targets for IACTs. Although blazars, particularly BL Lacs, dominate the extragalactic $\gamma$-ray sky, only a few dozen EHSPs are known (Nievas Rosillo et al., 2022; Acciari et al., 2020).
@@
The paper is organised as follows: Section 2 describes source selection and multi-wavelength data, Section 3 details SED modelling, and Section 4 discusses EHSP properties. Section 5 presents detectability predictions with Cherenkov Telescope Array Observatory (CTAO), and Section 6 provides a summary and conclusions.
@@ 2.1 Source selection
We select blazars from the Second Brazil-ICRANet Gamma-ray Blazars catalogue (2BIGB catalogue; Arsioli et al., 2020), which contains 1160 $\gamma$-ray blazars along with their $\gamma$-ray SEDs. These SEDs were obtained from the first 11 years of observations by the Large Area Telescope (LAT) aboard the Fermi Gamma-ray Space Telescope, a space-based instrument designed to detect $\gamma$ rays in the energy range from approximately 20 MeV to over 300 GeV (Atwood et al., 2009). The 2BIGB catalogue was produced through a $\gamma$-ray likelihood analysis of all sources in the 3HSP catalogue (Chang et al., 2019), which remains the largest collection of HSPs, EHSPs, and blazar candidates of uncertain type (BCUs, that is, sources detected by Fermi-LAT that show characteristics typical of blazars, as a double peaked SED, but lack definitive classification as either BL Lac or FSRQ).
To avoid contamination of diffuse emission from the Galactic plane, we select only blazars located at high latitudes ($|b|>10^{\circ}$). We also require that all sources in our sample have a redshift estimate, even if photometric, along with flux measurements in the optical/UV, X-ray, and $\gamma$-ray bands. With these criteria, we obtain a preliminary sample of 657 $\gamma$-ray blazars. We cross-reference these sources with the Fermi-LAT Fourth Source catalogue Data Release 4 (4FGL-DR4, Ballet et al., 2023), which lists $\gamma$-ray sources detected by Fermi-LAT over its first 14 years.
@@
Since we use archival data from multiple telescopes to reconstruct the broadband SEDs of the selected blazars, and these data are not always taken simultaneously, additional cuts are applied to select sources with low variability across different energy bands. The variability index from the 4FGL-DR4 catalogue is used to identify sources with low variability in the Fermi-LAT energy range. This index is calculated for each source as twice the sum of the log-likelihood differences between the flux in various time intervals and the average flux over the entire catalogue interval (Ajello et al., 2020). Only sources with a variability index below 27.69, the 99% confidence threshold in 4FGL-DR4, are included.
@@
To assess the variability of the sources in the X-ray band, data from the X-ray Telescope aboard the Neil Gehrels Swift Observatory (Swift-XRT, Burrows et al., 2005) are analysed to construct light curves. A Bayesian block analysis (Scargle et al., 2013), as implemented in astropy (Astropy Collaboration et al., 2022, 2018, 2013), is used to distinguish sources with a single block (likely non-variable) from those with multiple blocks (variable). We adopt a false probability $p_{0}=0.01$ and a systematic uncertainty of 10%. Only sources with a single Bayesian block in their light curve are included, resulting in 124 selected blazars. However, gaps in the temporal sampling may affect variability estimates. In general, variability studies benefit from frequent and continuous monitoring, which allows for better characterization of the synchrotron component. However, for EHSPs, the optical/UV bands often contain a significant contribution from the host galaxy’s thermal emission (e.g., Scarpa et al., 2000), reducing the sensitivity to AGN variability. As a result, no variability-based selection criteria are applied in this energy range. A skymap showing the distribution of the 124 selected blazars is presented in Figure 1.
@@
We do not apply any additional requirements regarding the class of the sources in 4FGL. Of the 124 sources in the final sample, 93 are classified as BL Lacs in the 4FGL-DR4 catalogue (Ballet et al., 2023), one as an FSRQ (4FGL J0132.7--0804, although it is likely misclassified; see Section 4.1), 1 as a radiogalaxy (4FGL J1518.6+0614), and 29 as BCUs. Most sources in the sample are BL Lacs since the blazars listed in the 3HSP catalogue show characteristics commonly associated with HSP and EHSP blazars. Figure 2 shows the dependence of the spectral index on flux and the variability index, as well as the relationship between the variability index and flux, for all sources in our sample and all sources classified as BL Lacs or FSRQs in the 4FGL catalogue. As expected, the sources selected for our study tend to have harder spectra than is typically found in FSRQs, as most of them belong to the BL Lac class.
@@
Figure 2: The 124 sources in our sample compared to all BL Lacs and FSRQs in the 4FGL-DR4 catalogue. Left: Photon spectral index versus the integral photon flux from 1 to 100 GeV. Middle: Photon spectral index versus the variability index. Right: Variability index versus the integral photon flux from 1 to 100 GeV. The photon index, flux, and variability index values are from the 4FGL-DR4 catalogue. We represent all the BL Lac sources from the 4FGL-DR4 catalogue with blue circles and the FSRQs with orange circles. The sources selected for this work are marked with dark crosses, the radiogalaxy of the sample (4FGL J1518.6+0614) is marked with a red star, and the FSRQ (4FGL J0132.7--0804) is marked with a green star.
The redshift distribution of the blazar sample is shown in Figure 3. They come from the 2BIGB (Arsioli et al., 2020) and 4LAC-DR3 (Ajello et al., 2020, 2022) catalogues, as well as from Goldoni (2021) and Paliya et al. (2021). Of these, 44 (35.5%) are from 4LAC, which does not specify whether they are spectroscopic or photometric, while 38 (30.6%) from 2BIGB are all photometric. The remaining redshifts, 33 (26.6%) from Paliya et al. (2021) and 9 (7.3%) from Goldoni (2021), are spectroscopic. Most of the sources in our sample ($\sim 86\%$) have a redshift of $z<0.5$, although some sources reach much higher redshifts, with the farthest blazar in the sample located at $z=2.075$ (4FGL J0323.7--0111, associated with 1RXS J032342.6--011131), though this redshift, given in 4LAC, may be uncertain as the optical spectrum of the associated source is continuum-dominated.
@@ 2.2 Multi-wavelength data
For the construction of the broadband SEDs of all sources in the sample, we gather data from various telescopes covering different energy bands. In the X-ray band, data from Swift-XRT are incorporated into the broadband SEDs of the sample blazars, using our own analysis of the data. We use the UKSSDC XRT-prods Python API [^1] https://www.swift.ac.uk/API (Evans et al., 2009, 2007), which retrieves high-level spectral data for the source and background regions along with the instrument response functions. Spectral parameters and flux points are then extracted with XSPEC (Arnaud, 1996). In the optical/UV range, we also include data from the UV/Optical Telescope aboard the Neil Gehrels Swift Observatory (Swift-UVOT, Roming et al., 2005), which we analyse using an automated tool built on the official HEASARC uvot-product package.
In the HE $\gamma$-ray band, we include flux data from Fermi-LAT available in the 4FGL-DR4 catalogue for eight energy bins. The only blazar in our sample detected in the VHE $\gamma$-ray band is 4FGL J0013.9--1854 (associated with SHBL J001355.9--185406), for which we also include the H.E.S.S. data provided in the Spectral TeV Extragalactic catalogue (STeVECat; Gréaux et al., 2023).
Additionally, for each source, we incorporate multi-wavelength archival data from the Space Science Data Center (SSDC) SED Builder service [^2] https://tools.ssdc.asi.it/SED/, an online tool for downloading data collected by multiple instruments over several decades. The SSDC SED Builder is used to extract historical data for each blazar across all bands except $\gamma$ rays. Since our study focuses on low-variability sources, we do not apply any time constraints on the SSDC data to exclude potential flaring periods.

### v2

@@ Abstract
Methods. A sample of 124 $\gamma$-ray blazars was analysed, selected for their high synchrotron peak frequencies and $\gamma$-ray emission properties, with a focus on sources showing low variability and good broadband data coverage. Their SEDs were constructed using archival multi-wavelength data from the SSDC SED Builder service, supplemented with recent Swift-UVOT, Swift-XRT, and Fermi-LAT observations. The SEDs were modelled with a one-zone synchrotron/synchrotron-self-Compton framework, classifying sources by synchrotron peak frequency. EHSP properties are compared to other blazar populations, and their detectability with the Cherenkov Telescope Array Observatory (CTAO) is assessed.
Results. We identify 66 new EHSP candidates, significantly expanding the known population. A clear correlation between synchrotron peak frequency and the magnetic-to-kinetic energy density ratio is found, with the most extreme EHSPs nearing equipartition. This indicates that as the synchrotron peak shifts to higher frequencies, the energy stored in the magnetic field becomes comparable to that of the relativistic electrons, suggesting a more balanced and energetically efficient jet environment in the most extreme blazars. Host galaxy emission is detected in many sources, but no significant differences are observed between elliptical and lenticular hosts. Finally, our analysis suggests that nine high-synchrotron peaked/EHSPs could be observed by CTAO at $>5\sigma$ (20 at $>3\sigma$) in 20-hour exposures, a feasible integration time for Imaging Atmospheric Cherenkov Telescopes, indicating that while the overall detection rate remains modest, a subset of these sources is within reach of next-generation very-high-energy gamma-ray instruments.
@@ 1 Introduction
Active galactic nuclei (AGNs) are extreme cosmic sources, powered by matter accretion onto a supermassive black hole (SMBH, e.g. Combes, 2021). Some develop relativistic jets extending to kiloparsec scales (e.g. Marschern, 2010). Unification models classify radio-loud AGNs by jet viewing angle (e.g. Urry & Padovani, 1995), with blazars, jets aligned close to our line of sight, dominating the extragalactic $\gamma$-ray sky (Abdollahi et al., 2020).
@@
Blazars have broadband spectral energy distributions (SEDs) dominated by non-thermal emission from radio to $\gamma$ rays (e.g. Fossati et al., 1998; Padovani, 2017). Their SEDs feature a double-peaked structure: the lower-energy peak (infrared to X-ray) arises from synchrotron emission by relativistic jet electrons, while the origin of the higher-energy $\gamma$-ray peak remains debated (e.g. Abdo et al., 2010), with both hadronic and leptonic processes proposed.
@@
Leptonic models explain the higher-energy peak of blazar SEDs as inverse Compton (IC) scattering, where high-energy electrons interact with low-energy photons (e.g. MAGIC Collaboration et al., 2020, 2021). This occurs mainly through synchrotron-self-Compton (SSC), where electrons scatter their own synchrotron photons (e.g. Maraschi et al., 1992; Finke et al., 2008), and external Compton, where electrons scatter external photons from sources like the cosmic microwave background, broad-line region (BLR), dusty torus, disc, stars, or different jet layers (e.g. Sikora et al., 1994; van den Berg et al., 2019). Hadronic models instead attribute $\gamma$-ray emission to high-energy proton interactions with ambient matter or radiation fields and to synchrotron radiation from relativistic protons in strong magnetic fields (e.g. Mannheim & Biermann, 1992; Böttcher et al., 2013, and references therein). Lepto-hadronic models combine both leptonic and hadronic processes (e.g. Mücke et al., 2003; Paliya et al., 2018). The main acceleration mechanisms proposed for ultra-relativistic particles in blazar jets include shock acceleration via repeated shock crossings and magnetic turbulence (e.g. Kirk et al., 1998; Tavecchio et al., 2020), as well as magnetic reconnection (e.g. Sironi & Spitkovsky, 2014) and stochastic acceleration (e.g. Blandford & Eichler, 1987; Liu et al., 2017).
@@
Blazars are classified based on their optical spectra into flat spectrum radio quasars (FSRQs) and BL Lacertae objects (BL Lacs). While FSRQs exhibit strong, broad emission lines (EW >5>5> 5 Å), BL Lacs have an almost featureless continuum (e.g. Marcha et al., 1996), indicating that external photon sources like the BLR, dusty torus, or disc are subdominant, with emission mainly from synchrotron/SSC (e.g. Ghisellini et al., 2011). In the most extreme BL Lacs, where the synchrotron peak is at higher frequencies, the host galaxy’s thermal emission becomes visible in the optical/ultraviolet (UV) range, allowing absorption lines to be detected. Blazars are also categorised by their synchrotron peak frequency (e.g. Padovani & Giommi, 1995; Abdo et al., 2010), from low-synchrotron peaked (LSP, $\nu_{sync}^{peak}<10^{14}$ Hz) to intermediate (ISP, $10^{14}$ Hz $\leq\nu_{sync}^{peak}<10^{15}$ Hz), high (HSP, $10^{15}$ Hz $\leq\nu_{sync}^{peak}<10^{17}$ Hz), and extreme high-synchrotron peaked (EHSP, $\nu_{sync}^{peak}\geq 10^{17}$ Hz), with peaks often reaching the X-ray band.
[…]
@@
The most luminous blazars, FSRQs, are mainly LSPs, while higher synchrotron peak frequency blazars are mostly lower-luminosity BL Lacs (e.g. Giommi et al., 2012). This anti-correlation between luminosity and synchrotron peak frequency defines the ‘blazar sequence’ (Fossati et al., 1998; Ghisellini et al., 2017), though observational biases and sample selection may challenge its validity (e.g. Padovani, 2007; Giommi et al., 2012; Padovani et al., 2022).
@@
Characterised by radiatively inefficient accretion discs, EHSPs occupy the end of the blazar sequence (e.g. Ghisellini et al., 2011). Their extreme synchrotron peak frequencies challenge standard models, requiring extreme conditions for the ultra-relativistic particles responsible for their emission (e.g. Biteau et al., 2020). Studying EHSPs enhances our understanding of jet acceleration and non-thermal emission. Their high energies (up to tens of tera-electronvolts, Biteau et al., 2020) make them valuable for cosmological studies, including the extragalactic background light (EBL, e.g. Dom´ınguez & Ajello, 2015; Dom´ınguez et al., 2019; Gréaux et al., 2024), blazar evolution (e.g. Ajello et al., 2014), the extragalactic $\gamma$-ray background (e.g. Ajello et al., 2015; Paliya et al., 2019), and the intergalactic magnetic field (e.g. Bonnoli et al., 2015). As potential very-high-energy (VHE, $E>100$ GeV) emitters, they are prime targets for IACTs. Although blazars, particularly BL Lacs, dominate the extragalactic $\gamma$-ray sky, only a few dozen EHSPs are known (Nievas Rosillo et al., 2022; Acciari et al., 2020).
@@
The paper is organised as follows: Section 2 describes source selection and multi-wavelength data, Section 3 details SED modelling, and Section 4 discusses EHSP properties. Section 5 presents detectability predictions with the Cherenkov Telescope Array Observatory (CTAO), and Section 6 provides a summary and conclusions.
@@ 2.1 Source selection
We selected blazars from the Second Brazil-ICRANet Gamma-ray Blazars catalogue (2BIGB catalogue; Arsioli et al., 2020), which contains 1160 $\gamma$-ray blazars along with their $\gamma$-ray SEDs. These SEDs were obtained from the first 11 years of observations by the Large Area Telescope (LAT) aboard the Fermi Gamma-ray Space Telescope, a space-based instrument designed to detect $\gamma$ rays in the energy range from approximately 20 MeV to over 300 GeV (Atwood et al., 2009). The 2BIGB catalogue was produced through a $\gamma$-ray likelihood analysis of all sources in the 3HSP catalogue (Chang et al., 2019), which remains the largest collection of HSPs, EHSPs, and blazar candidates of uncertain type (BCUs; that is, sources detected by Fermi-LAT that show characteristics typical of blazars, as a double peaked SED, but lack definitive classification as either BL Lac or FSRQ).
To avoid contamination of diffuse emission from the Galactic plane, we selected only blazars located at high latitudes ($|b|>10^{\circ}$). We also required that all sources in our sample have a redshift estimate, even if photometric, along with flux measurements in the optical/UV, X-ray, and $\gamma$-ray bands. With these criteria, we obtained a preliminary sample of 657 $\gamma$-ray blazars. We cross-referenced these sources with the Fermi-LAT Fourth Source catalogue Data Release 4 (4FGL-DR4, Ballet et al., 2023), which lists $\gamma$-ray sources detected by Fermi-LAT over its first 14 years.
@@
Since we used archival data from multiple telescopes to reconstruct the broadband SEDs of the selected blazars, and these data are not always taken simultaneously, additional cuts were applied to select sources with low variability across different energy bands. The variability index from the 4FGL-DR4 catalogue was used to identify sources with low variability in the Fermi-LAT energy range. This index is calculated for each source as twice the sum of the log-likelihood differences between the flux in various time intervals and the average flux over the entire catalogue interval (Ajello et al., 2020). Only sources with a variability index below 27.69, the 99% confidence threshold in 4FGL-DR4, were included.
@@
To assess the variability of the sources in the X-ray band, data from the X-ray Telescope aboard the Neil Gehrels Swift Observatory (Swift-XRT, Burrows et al., 2005) were analysed to construct light curves. A Bayesian block analysis (Scargle et al., 2013), as implemented in astropy (Astropy Collaboration et al., 2022, 2018, 2013), was used to distinguish sources with a single block (likely non-variable) from those with multiple blocks (variable). We adopted a false probability $p_{0}=0.01$ and a systematic uncertainty of 10%. Only sources with a single Bayesian block in their light curve were included, resulting in 124 selected blazars. However, gaps in the temporal sampling may affect variability estimates. In general, variability studies benefit from frequent and continuous monitoring, which allows for better characterisation of the synchrotron component. However, for EHSPs, the optical/UV bands often contain a significant contribution from the host galaxy’s thermal emission (e.g. Scarpa et al., 2000), reducing the sensitivity to AGN variability. As a result, no variability-based selection criteria were applied in this energy range. A sky-map showing the distribution of the 124 selected blazars is presented in Figure 1.
@@
We did not apply any additional requirements regarding the class of the sources in 4FGL. Of the 124 sources in the final sample, 93 are classified as BL Lacs in the 4FGL-DR4 catalogue (Ballet et al., 2023), one as an FSRQ (4FGL J0132.7--0804, although it is likely misclassified; see Section 4.1), one as a radiogalaxy (4FGL J1518.6+0614), and 29 as BCUs. Most sources in the sample are BL Lacs since the blazars listed in the 3HSP catalogue show characteristics commonly associated with HSP and EHSP blazars. Figure 2 shows the dependence of the spectral index on flux and the variability index, as well as the relationship between the variability index and flux, for all sources in our sample and all sources classified as BL Lacs or FSRQs in the 4FGL catalogue. As expected, the sources selected for our study tend to have harder spectra than is typically found in FSRQs, as most of them belong to the BL Lac class.
@@
Figure 2: Comparison of the 124 sources in our sample with all BL Lacs and FSRQs in the 4FGL-DR4 catalogue. Left: Photon spectral index versus the integral photon flux from 1 to 100 GeV. Middle: Photon spectral index versus the variability index. Right: Variability index versus the integral photon flux from 1 to 100 GeV. The photon index, flux, and variability index values are from the 4FGL-DR4 catalogue. We represent all the BL Lac sources from the 4FGL-DR4 catalogue with blue circles and the FSRQs with orange circles. The sources selected for this work are marked with dark crosses, the radiogalaxy of the sample (4FGL J1518.6+0614) is marked with a red star, and the FSRQ (4FGL J0132.7--0804) is marked with a green star.
[…]
@@ 2.2 Multi-wavelength data
For the construction of the broadband SEDs of all sources in the sample, we gathered data from various telescopes covering different energy bands. In the X-ray band, data from Swift-XRT were incorporated into the broadband SEDs of the sample blazars, using our own analysis of the data. We used the UKSSDC XRT-prods Python API [^1] https://www.swift.ac.uk/API (Evans et al., 2009, 2007), which retrieves high-level spectral data for the source and background regions along with the instrument response functions. Spectral parameters and flux points were then extracted with XSPEC (Arnaud, 1996). In the optical/UV range, we also included data from the UV/Optical Telescope aboard the Neil Gehrels Swift Observatory (Swift-UVOT, Roming et al., 2005), which we analysed using an automated tool built on the official HEASARC uvot-product package.
In the HE $\gamma$-ray band, we included flux data from Fermi-LAT available in the 4FGL-DR4 catalogue for eight energy bins. The only blazar in our sample detected in the VHE $\gamma$-ray band is 4FGL J0013.9--1854 (associated with SHBL J001355.9--185406), for which we also included the H.E.S.S. data provided in the Spectral TeV Extragalactic catalogue (STeVECat; Gréaux et al., 2023).
Additionally, for each source, we incorporated multi-wavelength archival data from the Space Science Data Center (SSDC) SED Builder service [^2] https://tools.ssdc.asi.it/SED/, an online tool for downloading data collected by multiple instruments over several decades. The SSDC SED Builder was used to extract historical data for each blazar across all bands except $\gamma$ rays. Since our study focuses on low-variability sources, we did not apply any time constraints on the SSDC data to exclude potential flaring periods.